OpenStudy (anonymous):

lim x-> sqrt(x-4)-1/x-13 HELP PLEASE

OpenStudy (anonymous):

\[\lim x->13 (\sqrt{x-4}-3)/x-13\]

OpenStudy (anonymous):

do you know L-Hospital rule??

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

if we get stuck in limit like this we apply L-Hospital rule which state.... take the derivative of numerator and denominator and then apply the given limit

OpenStudy (anonymous):

can you do it now??

OpenStudy (anonymous):

I'm not sure what you mean...ugh

OpenStudy (anonymous):

I come up with zero

OpenStudy (anonymous):

yes this is the right answer

OpenStudy (anonymous):

\[\lim_{x\to13}\frac{\sqrt{x-4}-3}{x-13}\cdot\frac{\sqrt{x-4}+3}{\sqrt{x-4}+3}\\ \lim_{x\to13}\frac{(x-4)-9}{(x-13)(\sqrt{x-4}+3)}\\ \lim_{x\to13}\frac{x-13}{(x-13)(\sqrt{x-4}+3)}\\ \lim_{x\to13}\frac{1}{\sqrt{x-4}+3}\]